0.7 pts Question 1 Fresh Foods is considering the purchase of a new packaging system. The system costs $83,732. The company pI NEED A CORRECT ANSWER PLS. PLS VERIFY THE ANSWER BEFORE POSTING.
Answer #1
Present Value of Cash Flow;
(Cash Flow)/((1+i)^N)
i=discount Rate=MARR=2%=0.02
N=Year of cash Flow
INTEREST AND PRINCIPAL REPAYMENT ON AMOUNT BORROWED
Pv Amount Borrowed =(3/4)*83732= $62,799
Nper Number of years of repayment                     6
Rate Interest Rate 8%
PMT Annual repayment for 6 years $13,584 (Using PMT function of excelwith Rate=8%,Nper=6, Pv=-62799
REPAYMENT SCHEDULE
Year 1 2 3 4 5 6
A Beginning Balance $62,799 $54,239 $44,993 $35,008 $24,225 $12,578
B Amount of annual payment $13,584 $13,584 $13,584 $13,584 $13,584 $13,584
C=A*8% Interest $5,024 $4,339 $3,599 $2,801 $1,938 $1,006
D=B-C Principal $8,560 $9,245 $9,985 $10,784 $11,646 $12,578
E=A-D Ending Balance $54,239 $44,993 $35,008 $24,225 $12,578 $0
Cash Invested in system =83732-62799 $20,933
20 $20.933 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 ($20,933) $3,738 -$5,024 -$850 $3,738 -$4,339 $850 $3,738 -$3,599 -$850 $3,H11 x fc =PMT(H10, H9,-H8) 7 8 Pv 9 Nper 10 Rate 11 PMT 12 H I J K L INTEREST AND PRINCIPAL REPAYMENT ON AMOUNT BORROWED AmouH33 : * ~ fic =PMT(2%,16,-432) к $3,738 $5,024 $850 $3,738 -$4,339 $850 $3,738 $3,599 $850 $3,738 $2,801 $850 $8,560 $9,245 $